OpenStudy (anonymous):

Let u = <-5, -9>, v = <6, 8>. Find -8u - 2v. please help :(

OpenStudy (amistre64):

its just multiplication and addition

OpenStudy (anonymous):

what do I multiply and add? Is there a formula?

OpenStudy (amistre64):

u = <-5, -9> v = < 6, 8> -8u = <-5(-8), -9(-8)> -2v = < 6(-2), 8(-2)> -8u = < 40, 72 > -2v = <-12, -16 > ------------------ ? , ?

OpenStudy (amistre64):

there is another notation if its simpler to see: u = <-5, -9> = -5x -9y v = < 6, 8> = 6x +8y -8u =-8(-5x -9y) -2v = -2(6x +8y) -8u = 40x + 72y -2v = -12x -16y -8u -2v = 40x + 72y -12x -16y and combine like terms

OpenStudy (anonymous):

Ahh okay I get it so my answer would be <28,56>??

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Thank you so much

OpenStudy (amistre64):

youre welcome